OpenStudy (anonymous):

Calculus III QUESTION. (a) Find the unit vectors that are parallel to the tangent line to the curve y = 8 sin x at the point (π/6, 4). (b) Find the unit vectors that are perpendicular to the tangent line.

OpenStudy (anonymous):

I don't know what I'm doing wrong I got, -1/2, (-1sqroot of3)/2 for a

OpenStudy (anonymous):

\[y=8\sin x~~\Rightarrow~~y'=8\cos x\] For \(x=\dfrac{\pi}{6}\), you have \(y'=8\cos\dfrac{\pi}{6}=\dfrac{8\sqrt3}{2}=4\sqrt3\). The tangent line would then be \[y-4=4\sqrt3\left(x-\frac{\pi}{6}\right)~~\iff~~y=4\sqrt3~x-\frac{2\sqrt3\pi}{3}+4\] Is this what you have?

OpenStudy (anonymous):

I found out what I did wrong. No that's not it. I got a different answer. y′ = 8cosx, so m = 8cos(π/6) = 4√3. with the vector ⟨1,4√3⟩ so you do the magnitude of that vector and you get 7. Then you divide the vector with the magnitude to get the unit vector, ⟨1/7,4√3/7⟩, then you also do the negative as well ⟨-1/7,-4√3/7⟩

OpenStudy (anonymous):

Thank God I got it right. Got a 100 percent!!!

OpenStudy (anonymous):

Nice job!

OpenStudy (anonymous):

Thanks. Thanks for trying to help though